# EEG-based Graph-guided Domain Adaptation for Robust Cross-Session Emotion Recognition

### v1

@@ V Conclusion
In this work, we proposed EGDAframework for robust cross-session emotion recognition. The proposed method integrates distribution alignment with graph regularization to simultaneously reduce marginal and conditional discrepancies across sessions. By minimizing within-class scatter matrices in the source domain, EGDAencourages compact and discriminative feature representations, thereby facilitating more effective domain adaptation. Experimental results on the SEED-IV dataset demonstrate that EGDAachieves superior recognition accuracy compared with several state-of-the-art baseline methods. Furthermore, our analysis highlights the Gamma frequency band as the most discriminative and identifies the prefrontal, central-parietal, and temporal brain regions as critical for emotional processing.

### v2

@@ V Conclusion
In this work, we proposed the EGDA framework for robust cross-session emotion recognition. The proposed method integrates distribution alignment with graph regularization to simultaneously reduce marginal and conditional discrepancies across sessions. By minimizing within-class scatter matrices in the source domain, EGDA encourages compact and discriminative feature representations, thereby facilitating more effective domain adaptation. Experimental results on the SEED-IV dataset demonstrate that EGDA achieves superior recognition accuracy compared with several state-of-the-art baseline methods. Furthermore, our analysis highlights the Gamma frequency band as the most discriminative and identifies the prefrontal, central-parietal, and temporal brain regions as critical for emotional processing.